# Fractal and Regular Geometry of Deep Neural Networks

### v1

@@ Abstract
We study the geometric properties of random neural networks by investigating the boundary volumes of their excursion sets for different activation functions, as the depth increases. More specifically, we show that, for activations which are not very regular (e.g., the Heaviside step function), the boundary volumes exhibit fractal behavior, with their Hausdorff dimension monotonically increasing with the depth. On the other hand, for activations which are more regular (e.g., ReLU, logistic and $\tanh$), as the depth increases, the expected boundary volumes can either converge to zero, remain constant or diverge exponentially, depending on a single spectral parameter which can be easily computed. Our theoretical results are confirmed in some numerical experiments based on Monte Carlo simulations.
@@ 1   Introduction
In recent years, neural networks have attracted considerable attention from mathematicians, leading to significant advances in their theoretical understanding. Efforts to analyze deep learning have drawn from various perspectives, including approximation theory [22], statistics [44, 9], and optimization [3, 18]. Beside these approaches, another line of research has centered on the study of geometric properties, with the goal of characterizing the complexity of neural networks across different architectures [42, 34, 32, 8, 11, 46]. Meanwhile, building upon a classical result by Neal [43], several papers have studied large-width networks at initialization through the lenses of Gaussian processes, proving central limit theorems at different levels of generality.
From this latter point of view, one considers a neural network with suitably normalized random weights, and let the width grow to infinity. Neal first observed that any evaluation of a shallow network converges to a Gaussian random variable. Later on, this was extended to deep networks seen as random functions, showing convergence to a Gaussian process, see the functional central limit theorems [21, 39, 33] and their quantitative counterparts [14, 28, 10]. As a consequence, random neural networks with large width are very close to Gaussian fields. In some sense, this could be even viewed as a characterization of neural networks for “generic” values of their parameters, although it is not guaranteed that the networks will show a similar behavior after training.
@@
A first step in this direction was taken in [24]. The authors considered isotropic Gaussian random fields on the sphere $\mathbb{S}^{d}$ emerging from random neural networks, and studied their angular power spectrum as a function of the nonlinear activation function $\sigma$ and the depth $L$ of the network. This explicit analysis led to a classification of networks into three classes:
@@
The present paper builds from that contribution to explore the geometric features of neural networks in the infinite-width limit. In particular, we investigate how the geometry of their random excursion sets is influenced by the architecture (activation function and depth). We define the Covariance Regularity Index (CRI) of a network (Definition 3.5), describing the regularity of the kernel of its associated Gaussian field; we also explore the connection that exists between the CRI and the decay of the angular spectrum, see Proposition 3.4). The value of the CRI naturally leads to the classification of random neural networks into two classes with drastically different geometric behavior:
@@ 1. Fractal class:
    A CRI value smaller than $1$, corresponding to activation functions that are not regular (e.g., the Heaviside activation), yields excursion volumes that show fractal behavior, that is, non-integer Hausdorff dimension (Theorem 3.9). More than that, the Hausdorff dimension of the excursion sets grows as the depth $L$ increases, converging to the dimension of the ambient space. In this sense, we can say that excursion sets of these random neural networks are “more and more fractal” for deeper and deeper architectures.
@@ 2. Kac-Rice class:
    Networks with a CRI value larger than $1$ always exhibit excursion volumes of finite size (Theorem 3.12), with an explicit finite expectation. The behavior of this expectation as the depth $L$ increases, though, is greatly influenced by the underlying activation function. Indeed, neural networks falling in this class are further split into three categories, corresponding to the three regimes of [24] (Remark 3.13): neural networks belonging to the low-disorder regime have excursion volumes converging to zero on average as $L$ grows to infinity; for neural networks in the high-disorder regime these expectations diverge exponentially; for the sparse regime (which includes the ReLU activation), the expected boundary volume of excursion sets is constant.
@@
For the fractal class, we carefully investigate the strong local nondeterministic properties of the field, see [20, 48, 49, 36] for an overview on this topic. We extend to any dimension a number of results for non-regular Gaussian fields from [40], where the authors only dealt with fields having domain on the $2$-dimensional sphere. In particular, we compute the exact uniform modulus of continuity for an isotropic Gaussian random field with CRI smaller than $1$ (Proposition 4.4); using this modulus, we upper bound the Hausdorff dimension of the graph of the field by an explicit construction of a $\delta$-net covering, and obtain a lower bound using a standard argument based on the potential method. For the dimension of the boundary excursion set, the lower bound is computed using again the potential method, while for the upper bound we use a generalization to the sphere of a classical result on the dimension of sets intersections with parallel planes. The proof is then completed by observing that the CRI of Gaussian fields resulting from neural networks in the fractal class decreases exponentially fast to 0 as L goes to infinity.
For the regular class, the idea is to use the classical Kac-Rice approach (see [2, 7]). In particular, we use the results from the recent paper [5] and combine them together with the characterization of the angular power spectrum derived in [12]. In order to apply the Kac-Rice machinery, though, one needs to check that the Gaussian field is $C^{1}$-smooth; this property is well known to be satisfied when the field has CRI equal to $2$ (that is, when its kernel is $C^{2}$ in the whole interval $[-1,1]$), but this leaves out fields induced by activation functions (including ReLU) yielding a CRI between $1$ and $2$. Thus we fill this gap by proving that any Gaussian field with CRI larger than $1$ is almost surely $C^{1}$, see Proposition 3.14. This result is obtained through a stereographic projection of the field and a thorough analysis of its mean-square derivatives, leading to an application of Kolmogorov’s continuity theorem.
Once we have established that the formulas for calculating the expected excursion volumes from  [5] can be applied to any neural network in the Kac-Rice class, we explicitly calculate these expectations, proving that their value only depends on the derivative of the kernel in a neighborhood of $1$. This value, in turn, is exactly what distinguished the three regimes appearing in [24].
@@ 2.1   Spherical random fields
[…]
@@
  • $n_{\ell,d}$ is the number of linearly indipendent homogeneous harmonic polynomials of degree $\ell$ in $d+1$ variables given by $n_{0,d}=1$ and for $\ell>0$
    $n_{\ell,d}$ $=\frac{2\ell+d-1}{\ell}\binom{\ell+d-2}{\ell-1}\ ;$
@@
  • $(G_{\ell;d})_{\ell\in\mathbb{N}}$ is a sequence of polynomials known as normalized Gegenbauer polynomials (see Section A.1 for more details).
@@
    where $\operatorname{dVol}$ denotes the $d$-dimensional volume form on $\mathbb{S}^{d}$. See Appendix B for an explicit construction. If $N_{d}=(0,\dotsc,0,1)\in\mathbb{S}^{d}$ denotes the north pole, one can choose the spherical harmonics such that
@@
    See Lemma B.1 for more details.
@@ 2.2   Random neural networks
For $L\geq 1$, let $n_{0}=d,n_{1},\dots,n_{L},n_{L+1}$ be positive integers. Let $\sigma:\mathbb{R}\to\mathbb{R}$ be a (nonlinear) function. A neural network of depth $L$, widths $n_{1},\dots,n_{L}$ and activation $\sigma$ is a function $T_{L}:\,\mathbb{S}^{d}\to\mathbb{R}^{n_{L+1}}$ defined by the following recursive formula:
@@
where $W^{(s)}\in\mathbb{R}^{n_{s+1}\times n_{s}}$ and $b^{(s)}\in\mathbb{R}^{n_{s}}$ are called weights and biases, respectively, and $\sigma$ is understood to be applied entrywise. We say that $T_{L}$ is a random neural network if weights and biases are independent random variables with independent components. For $\Gamma_{b}\in[0,1)$ we take the following standard calibration condition:
@@ 3.1   Spectral index and activation functions
Let $T$ be a random field and let $\kappa$ be its covariance function. Assume that $\kappa\in C^{2}((-1,1))$.
  • We say that $T$ has Covariance Regularity Index (CRI) equal to $2$ if $\kappa\in C^{2}([-1,1])$.
  • We say that $T$ has CRI equal to $\beta$, with $\beta<2$, if $\kappa$ admits the following expansions around ±$\pm 1$ as $t\to 0^{+}$:
@@
    where $p_{1},p_{-1}$ are polynomials and $c_{1},c_{-1}$ are constants, and the first two derivatives of $\kappa$ admit similar expansions obtained by differentiating the above ones.
Note that, if $\kappa$ is $\Delta$-analytic, then the last condition on the derivatives of $\kappa$ can be omitted.
We want to establish a connection between the decay of the angular powers spectrum of $T$ and its CRI. To this end, we introduce the concept of spectral index.
@@
where $q(z)>0$ is a polynomial. Then we say that $T$ has spectral index $\alpha$.
@@
The following proposition links the two definitions and it is proved in Section A.2.
@@ 3.2   The fractal class
Our first results cover the cases where the activation function is not regular enough to ensure that the excursion sets have integer Hausdorff dimension; in other words, we are here in the fractal regime. Throughout the paper, we will denote by $\dim_{\rm H}$ the Hausdorff dimension (see Section 4.1 for more details).
@@
Let $\sigma$ be an activation function in the fractal class. For any $L$, let $T_{L}$ be a random neural network with activation $\sigma$ and depth $L$. If $\beta\in(0,1)$ is the CRI of $T_{1}$, then the CRI of $T_{L}$ is $\beta^{L}$. In particular, the CRI is in $(0,1)$ for any $L$. Moreover,
@@
  • for almost all $u$ (in the sense of $1$-Lebesgue measure) and with probability one,
    $\dim_{\rm H}(T_{L}^{-1}(u))=d-\beta^{L}\ .$ (3.3)
@@
Using classical results (see e.g. [19]), one can also prove that, for any given $u$, we have
$\dim_{\rm H}(T_{L}^{-1}(u))=d-\beta^{L}$
with positive probability. Hence, for all $u\in\mathbb{R}$,
$\mathbb{E}\big{[}\mathcal{H}^{d-1}(A_{u}(T_{L}))\big{]}=+\infty\ .$
@@
Note that $\beta^{L}\to 0$ as $L\to+\infty$, and the Hausdorff dimension of $A_{u}$ converges to the Hausdorff dimension of the ambient space.
@@ 3.3   The Kac-Rice class
In this subsection we study the case where the random neural networks are regular enough to exhibit standard, integer-valued Hausdorff dimension of the boundary volumes (namely $\dim_{\rm H}(A_{u})=d-1$). On the other hand, the expected values of these volumes depend heavily on the depth of the networks, as illustrated by Theorem 3.12.
@@
$\mathbb{E}\left[\mathcal{H}^{d-1}(A_{u}(T_{L}))\right]=\omega_{d-1}\kappa^{\prime}(1)^{L/2}e^{-u^{2}/2}$
@@
This result mirrors the partition of neural networks into three classes (low-disorder, sparse and high-disorder) established in the recent paper [24]. In particular, in the low-disorder case ($\kappa^{\prime}(1)<1$) the expected boundary volumes converge to 00 as $L\to\infty$, in the sparse case ($\kappa^{\prime}(1)=1$) they remain constant on average, and in the high-disorder case ($\kappa^{\prime}(1)>1$) their expected values diverge exponentially fast.
To prove the previous results we need that $T_{L}\in C^{1}(\mathbb{S}^{d})$ a.s.; this is well-known to be satisfied when $\kappa\in C^{2}([-1,1])$ (i.e. $T_{L}$ has CRI equals to $2$) (see e.g. [38]). However, this leaves out important cases, in particular the celebrated ReLU activation function. Our next result shows that, if $\sigma$ belongs to the Kac-Rice class, then $T_{L}\in C^{1}(\mathbb{S}^{d})$ a.s.. This result has independent interest for the theory of geometry of random fields, because it allows to exploit the Kac-Rice approach in more general settings than the ones given for instance in [2, 7].
@@
4   Proof of Theorem 3.9: fractal class

### v2

@@ Abstract
We study the geometric properties of random neural networks by investigating the boundary volumes of their excursion sets for different activation functions, as the depth increases. More specifically, we show that, for activations which are not very regular (e.g., the Heaviside step function), the boundary volumes exhibit fractal behaviour, with their Hausdorff dimension monotonically increasing with the depth. On the other hand, for activations which are more regular (e.g., ReLU, logistic and $\tanh$), as the depth increases, the expected boundary volumes can either converge to zero, remain constant or diverge exponentially, depending on a single spectral parameter which can be easily computed. Our theoretical results are confirmed in some numerical experiments based on Monte Carlo simulations.
@@ 1  Introduction
In recent years, neural networks have attracted considerable attention from mathematicians, leading to significant advances in their theoretical understanding. Efforts to analyse deep learning have drawn from various perspectives, including approximation theory [24], statistics [49, 11], and optimization [4, 21]. Besides these approaches, another line of research has centred on the study of geometric properties, with the goal of characterizing the complexity of neural networks across different architectures [47, 37, 35, 10, 13, 51]. Meanwhile, building upon a classical result by Neal [48], several papers have studied large-width networks at initialization through the lenses of Gaussian processes, proving central limit theorems at different levels of generality.
From this latter point of view, one considers a neural network with suitably normalized random weights, and let the width grow to infinity. Neal first observed that any evaluation of a shallow network converges to a Gaussian random variable. Later on, this was extended to deep networks seen as random functions, showing convergence to a Gaussian process, see the functional central limit theorems [23, 43, 36] and their quantitative counterparts [16, 12, 31, 18], as well as large deviation properties [44, 5, 55]. As a consequence, random neural networks with large width are very close to Gaussian fields. In some sense, this could be even viewed as a characterization of neural networks for “generic” values of their parameters, although it is not guaranteed that the networks will show a similar behaviour after training.
@@
A first step in this direction was taken in [27]. The authors considered isotropic Gaussian random fields on the sphere $\mathbb{S}^{d}$ emerging from random neural networks, and studied their angular power spectrum as a function of the non-linear activation function $\sigma$ and the depth $L$ of the network. This explicit analysis led to a classification of networks into three classes:
@@
The present paper builds from that contribution to explore the geometric features of neural networks in the infinite-width limit. In particular, we investigate how the geometry of their random excursion sets is influenced by the architecture (activation function and depth). We define the Covariance Regularity Index (CRI) of a network (Definition˜3.6), describing the regularity of the kernel of its associated Gaussian field; we also explore the connection that exists between the CRI and the decay of the angular spectrum, see Proposition˜3.4). The value of the CRI naturally leads to the classification of random neural networks into two classes with drastically different geometric behaviour:
@@ 1. Fractal class:
    A CRI value smaller than $1$, corresponding to activation functions that are not regular (e.g., the Heaviside activation), yields excursion volumes that show fractal behaviour, that is, non-integer Hausdorff dimension (Theorem˜3.11). More than that, the Hausdorff dimension of the excursion sets grows as the depth $L$ increases, converging to the dimension of the ambient space. In this sense, we can say that excursion sets of these random neural networks are “more and more fractal” for deeper and deeper architectures.
@@ 2. Kac-Rice class:
    Networks with a CRI value larger than $1$ always exhibit excursion volumes of finite size (Theorem˜3.14), with an explicit finite expectation. The behaviour of this expectation as the depth $L$ increases, though, is greatly influenced by the underlying activation function. Indeed, neural networks falling in this class are further split into three categories, corresponding to the three regimes of [27] (Remark˜3.15): neural networks belonging to the low-disorder regime have excursion volumes converging to zero on average as $L$ grows to infinity; for neural networks in the high-disorder regime these expectations diverge exponentially; for the sparse regime (which includes the ReLU activation), the expected boundary volume of excursion sets is constant.
@@
For the fractal class, we carefully investigate the strong local nondeterministic properties of the field, see [22, 53, 54, 39] for an overview on this topic. We extend to any dimension a number of results for non-regular Gaussian fields from [45], where the authors only dealt with fields having domain on the $2$-dimensional sphere. In particular, we compute the exact uniform modulus of continuity for an isotropic Gaussian random field with CRI smaller than $1$ (Proposition˜4.5); using this modulus, we upper bound the Hausdorff dimension of the graph of the field by an explicit construction of a $\delta$-net covering, and obtain a lower bound using a standard argument based on the potential method. For the dimension of the boundary excursion set, the lower bound is computed using again the potential method, while for the upper bound we use a generalization to the sphere of a classical result on the dimension of sets intersections with parallel planes. The proof is then completed by observing that the CRI of Gaussian fields resulting from neural networks in the fractal class decreases exponentially fast to 0 as L goes to infinity.
For the regular class, the idea is to use the classical Kac-Rice approach (see [3, 9]). In particular, we use the results from the recent paper [7] and combine them together with the characterization of the angular power spectrum derived in [14]. In order to apply the Kac-Rice machinery, though, one needs to check that the Gaussian field is $C^{1}$-smooth; this property is well known to be satisfied when the field has CRI equal to $2$ (that is, when its kernel is $C^{2}$ in the whole interval [1,1][-1,1]), but this leaves out fields induced by activation functions (including ReLU) yielding a CRI between $1$ and $2$. Thus we fill this gap by proving that any Gaussian field with CRI larger than $1$ is almost surely $C^{1}$, see Proposition˜3.16. This result is obtained through a stereographic projection of the field and a thorough analysis of its mean-square derivatives, leading to an application of Kolmogorov’s continuity theorem.
Once we have established that the formulas for calculating the expected excursion volumes from  [7] can be applied to any neural network in the Kac-Rice class, we explicitly calculate these expectations, proving that their value only depends on the derivative of the kernel in a neighbourhood of $1$. This value, in turn, is exactly what distinguished the three regimes appearing in [27].
The plan of the paper is as follows: in Section˜2 we provide background material and the definitions that we use throughout the paper.  Section˜3 contains all our main results, including fractal behaviour and finite volumes of the excursion boundary. We show the main steps of the proof for the fractal case in Section˜4, and for the finite volumes case in Section˜6. More technical results and auxiliary lemmas are collected in the Appendix. In Section˜7 we illustrate our results by numerical simulations in dimension $2$.
@@ 2.1 Spherical random fields
where $\langle\cdot,\cdot\rangle$ denotes the scalar product in $\mathbb{R}^{d+1}$.
Hence, by Schoenberg’s theorem [50], the covariance can be expressed as
@@
  • $n_{\ell,d}$ is the number of linearly independent homogeneous harmonic polynomials of degree $\ell$ in $d+1$ variables given by $n_{0,d}=1$ and for $\ell>0$
    $n_{\ell,d}$ $=\frac{2\ell+d-1}{\ell}\binom{\ell+d-2}{\ell-1}\$;
@@
  • $(G_{\ell;d})_{\ell\in\mathbb{N}}$ is a sequence of polynomials known as normalized Gegenbauer polynomials (see Section˜A.1 for more details).
@@
    where $\operatorname{dVol}$ denotes the $d$-dimensional volume form on $\mathbb{S}^{d}$. See Appendix˜B for an explicit construction. If $N_{d}=(0,\dotsc,0,1)\in\mathbb{S}^{d}$ denotes the north pole, one can choose the spherical harmonics such that
@@
    See Lemma˜B.1 for more details.
@@ 2.2 Random neural networks
For $L\geq 1$, let $n_{0}=d,n_{1},\dots,n_{L},n_{L+1}$ be positive integers. Let $\sigma:\mathbb{R}\to\mathbb{R}$ be a (non-linear) function. A neural network of depth $L$, widths $n_{1},\dots,n_{L}$ and activation $\sigma$ is a function $T_{L}:\,\mathbb{S}^{d}\to\mathbb{R}^{n_{L+1}}$ defined by the following recursive formula:
@@
where $W^{(s)}\in\mathbb{R}^{n_{s+1}\times n_{s}}$ and $b^{(s)}\in\mathbb{R}^{n_{s}}$ are called weights and biases, respectively, and $\sigma$ is understood to be applied entry-wise. We say that $T_{L}$ is a random neural network if weights and biases are independent random variables with independent components. For $\Gamma_{b}\in[0,1)$ we take the following standard calibration condition:
@@ 3.1 Spectral index and activation functions
Let $T$ be a random field and let $\kappa$ be its covariance function. Assume that $\kappa\in C^{2}((-1,1))$. The CRI of $T$ is the largest $\beta\in(0,2]$ such that
  • $\kappa$ admits the following expansions around $\pm 1$ as $t\to 0^{+}$:
@@
    where $p_{1},p_{-1}$ are polynomials of degree strictly smaller than $\beta$, and $c_{1},c_{-1}\neq 0$ are constants;
  • $\kappa^{\prime}$ and $\kappa^{\prime\prime}$ admit similar expansions obtained by differentiating (3.1).
Note that, if $\kappa$ is $\Delta$-analytic, then the last condition on the derivatives of $\kappa$ can be omitted. We also remark that the CRI is equal to $2$ if and only if $\kappa\in C^{2}([-1,1])$. More in general, a CRI equal to $\beta$ implies that $\kappa$ is Hölder continuous of order $\beta$. It follows that a higher CRI of the random field implies higher Hölder regularity of the covariance function, and thus of the field itself (see Proposition˜3.16 and [25, Proposition 3.3] for a generalization).
We want to establish a connection between the decay of the angular power spectrum of $T$ and its CRI. To this end, we introduce the concept of spectral index.
@@
where $q$ is a polynomial such that $q(z)>0$ for all $z\in\mathbb{R}$. Then we say that $T$ has spectral index $\alpha$.
@@
The following proposition links the two definitions and it is proved in Section˜A.2.
@@
Remark 3.5.
Using the computation in Section˜A.2, for $\alpha=2$ we have
$\kappa(1-t)=1+Bt\ln(t)+o(t\ln(t))\ ,$
that is, there is a logarithmic correction. Hence Proposition˜3.4 does not hold for $\alpha=2$.
@@
The critical case CRI equal to $1$ is not included in any of the two classes above because, as we will show in Theorem˜3.11, the excursion sets are not fractal, nor the field is sufficiently regular to use the Kac-Rice formula.
@@ 3.2 The fractal class
Our first result covers the case where the activation function is not regular enough to ensure that the excursion sets have integer Hausdorff dimension, i.e. the fractal regime. Throughout the paper, we will denote by $\dim_{\rm H}$ the Hausdorff dimension (see Section˜4.1 for more details).
@@
Let $\sigma$ be an activation function in the fractal class. For any $L$, let $T_{L}$ be a random neural network with activation $\sigma$ and depth $L$. If $\beta\in(0,1]$ is the CRI of $T_{1}$, then the CRI of $T_{L}$ is $\beta^{L}$. In particular, the CRI is in $(0,1]$ for any $L$. Moreover,
@@
  • for every $u\in\mathbb{R}$,
    $\dim_{\rm H}(T_{L}^{-1}(u))=d-\beta^{L}$ (3.3)
    with positive probability.
@@
Theorem˜3.11 implies that for $\beta\in(0,1)$ and for all $u\in\mathbb{R}$,
$\mathbb{E}\big[\mathcal{H}^{d-1}(T_{L}^{-1}(u))\big]=+\infty\ .$
@@
Note that, if $\beta\in(0,1)$, then $\beta^{L}\to 0$ as $L\to+\infty$, and the Hausdorff dimension of $T_{L}^{-1}(u)$ converges to the Hausdorff dimension of the ambient space.
@@ 3.3 The Kac-Rice class
In this subsection, we study the case where the random neural network is regular enough to exhibit standard, integer-valued Hausdorff dimension of the boundary volumes (namely $\dim_{\rm H}(T_{L}^{-1}(u))=d-1$). On the other hand, the expected values of these volumes depend heavily on the depth of the network, as illustrated by Theorem˜3.14.
@@
$\mathbb{E}\left[\mathcal{H}^{d-1}(T_{L}^{-1}(u))\right]=\omega_{d-1}\kappa^{\prime}(1)^{L/2}e^{-u^{2}/2}$
@@
In the recent paper [27], neural networks were divided into three classes with different $L^{2}$ convergence regimes: low-disorder ($\kappa^{\prime}(1)<1$), sparse ($\kappa^{\prime}(1)=1$) and high-disorder ($\kappa^{\prime}(1)>1$). Theorem˜3.14 confirms geometrically what [27] established analytically. In particular, in the low-disorder case, when the field converges (in $L^{2}$) to a constant as $L\to\infty$, the expected boundary volumes converge to 0. In the sparse case, when the field converges to a constant but its derivatives do not vanish, the boundary volumes remain constant on average. In the high-disorder case, when the field does not converge to a constant and its derivatives diverge exponentially, the expected boundary volumes also diverge exponentially fast.
To prove the previous results we need that $T_{L}\in C^{1}(\mathbb{S}^{d})$ a.s.; this is well-known to be satisfied when $\kappa\in C^{2}([-1,1])$ (i.e. $T_{L}$ has CRI equal to $2$) (see e.g. [42]). However, this leaves out important cases, in particular the celebrated ReLU activation function. Our next result shows that, if $\sigma$ belongs to the Kac-Rice class, then $T_{L}\in C^{1}(\mathbb{S}^{d})$ a.s.. This result has independent interest for the theory of the geometry of random fields, because it allows us to exploit the Kac-Rice approach in more general settings than the ones given for instance in [3, 9].
@@
4  Proof of Theorem˜3.11: fractal class